How do I simplify the expression -5+i/2i
Answer by stanbon(75887)

How do I simplify the expression -5+i/2i
-------------
Since you did not use parentheses, the problem
can be interpreted in more than one way::
1) -5 + i/(2i) = -5 + (1/2) = -4.5
---------------
2) -5 + (i/2)i = -5 + (-1/2) = -5.5
-----------------
3) (-5+i)/(2i) = (-5/(2i) + (i/(2i) = (-5i/(-2) + 1/2) = (1/2) +(5/2)i
-------------------------------------------------------
Note:: Avoid ambiguity--- use parentheses please.
Cheers,
Stan H.